Part A Constants A man pushes on a piano with mass 170 kg ;it slides at constant velocity down a ramp that is inclined at 12.0° above the horizontal floor. Neglect any friction acting on the piano. Calculate the magnitude of the force applied by the man if he pushes parallel to the incline. Express your answer with the appropriate units. F1083 Submit X Incorrect; Try Again; 4 attempts remaining Part B Calculate the magnitude of the force applied by the man if he pushes parallel to the floor. Express your answer with the appropriate units. F-455 Submit
Answer #1

part A

component of force acting down the plane = Fg = m*g*sin\theta

let F be the force applied on the piano up the plane parallel to incline

Fnet = m*g*sin\theta - F

the piano is moving with constant velocity

acceleration = a = 0

Fnet = m*a = 0

m*g*sin\theta = F

F = 170*9.81*sin12

F = 346.73 N

part B

component of force paralle to the incline = F*cos12

F*cos12 = m*g*sin12

F = m*g*tan12

F = 354.11 N